# Efficiently Computing Equilibria in Budget-Aggregation Games

### v2

@@ Abstract
Budget aggregation deals with the social choice problem of distributing an exogenously given budget among a set of public projects, given agents’ preferences. Taking a game-theoretic perspective, we initialize the study of budget-aggregation games where each agent has virtual decision power over some fraction of the budget. This paper investigates the structure and shows efficient computability of Nash equilibria in this setting for various preference models. In particular, we show that Nash equilibria for Leontief utilities can be found in polynomial time, solving an open problem from Brandt et al. (2023).
@@ 1 Introduction
Participatory budgeting (see, e.g., Cabannes, 2004) is a democratic process in which citizens are allowed to vote on how the funds of their city or state should be distributed. It is justified both by ethical arguments (people have a fundamental right to decide how their money is used) and by practical arguments (letting people decide on the budget strengthens cohesion in the community).
This approach can be taken to the extreme: instead of just letting citizens vote on which projects they approve (as is commonly done today), one can give each citizen their own (virtual) share of the budget (e.g. $1/n$ of the total budget, in case there are $n$ citizens with equal rights), and allow them to decide how to allocate that share among the potential projects. This can be an attractive approach, as it gives full power to the citizens and ensures some degree of fairness by splitting this power equally among them. Furthermore, each citizen is able to directly observe her influence on the outcome by means of the distribution of her individual budget.
However, this approach raises severe strategic problems. For example, if a citizen thinks that projects $x$ and $y$ should each receive $0.5$ of the total budget, but later observes that most other citizens allocated their shares to $x$, then she would have preferred to spend her entire budget on $y$, leaving her unsatisfied since her budget was spent suboptimally in hindsight. This bad user experience can be avoided if we have the possibility to coordinate individual budgets by advising each citizen in advance how to distribute their own budget such that for each citizen, the advised distribution is optimal, given that all other citizens follow the provided advice.
@@
We propose a new class of normal-form games, called budget-aggregation games, which model (equal-resource) public funding situations. We position these games within the broader framework of normal-form games and survey existing results on the computation of NEs. On top of that, our paper analyzes NEs for the most common utility functions that have been considered in participatory budgeting and related fields. We prove theorems on the structure of NEs, and use these theorems to develop algorithms to efficiently compute them. We focus on the following utility models:
@@
  • For Leontief utilities, we show that an NE can be computed in polynomial time using a variant of the ellipsoid method, solving an open problem posed in Brandt et al. (2023) (Section 5).
@@
Finally, we extend the basic model by allowing agents to have different weights. We show that, for most of the above utility models, a NE can be computed in pseudo-polynomial time (Section 8).
@@ 2 Related Work
An important generalization of Nash’s theorem involves the notion of a concave game (Rosen, 1965). A concave game is defined by, for each player $i$, a convex set $(S_{i})_{i\in N}$ that represents the possible strategies for $i$, and a utility function $u_{i}$ that maps each strategy profile to a real number; $u_{i}$ can be any function that is continuous in all strategies, and concave in the strategy of ii. [^1] The definition in Rosen (1965) is even more general and allows coupled constraints, that is, the set of strategies available to each agent may depend on the choices made by other agents. We do not need this added generality here. Nash’s game is a special case in which every $S_{i}$ is a standard simplex in some Euclidean space (representing the set of all lotteries over pure actions), and every $u_{i}$ is a linear combination of products of $n$ terms. Our budget-aggregation game is a special case in which every $S_{i}$ is (still) a simplex, but the utility functions are substantially different, due to the interpretation as a distribution of some exogenously given budget rather than a lottery.
2.1 Computing an equilibrium is hard in general
Like most fixed-point-based proofs, the proofs of Nash (1950, 1951) and Rosen (1965) do not yield efficient algorithms for finding an equilibrium. Moreover, there are 3-player games with integer payoffs with a unique (mixed) NE in which all probabilities are irrational numbers (Nash, 1951; Bilò and Mavronicolas, 2014), and 4-player games with a unique NE in which all probabilities are irradical numbers (cannot be expressed using integer roots) (Orzech and Rinard, 2025). This shows that no finite algorithm can compute an exact NE in general games.
Even computing an $\epsilon$-approximate NE probably cannot be done in polynomial time, as it is PPAD-hard even for two players. This was first proven for $\epsilon$ that shrinks exponentially in the number of actions (Chen and Deng, 2006), then for $\epsilon$ that shrinks polynomially (Chen et al., 2006), and later even for some constant $\epsilon>0$ (Rubinstein, 2015). Computing an equilibrium is PPAD-complete, also for some other classes of concave games (Papadimitriou et al., 2023).
There are other classes of games for which an efficient algorithm is known. We survey some of them below.
@@ 2.2 Nash games
For the original game studied by Nash (where the strategy set of each agent is the set of lotteries over a finite set of actions), many practical algorithms for computing an NE were designed, such as the ones by Lemke and Howson (1964) and Porter et al. (2008), but they do not have a polynomial-time guarantee.
Polynomial-time algorithms for $\epsilon$-approximate NE are known for $\epsilon\approx 0.34$ for two players (Tsaknakis and Spirakis, 2007), $\epsilon\approx 0.5$ for $n$-player polymatrix games, and $\epsilon>0.6$ for general $n$-player games with $n\geq 3$ (Deligkas et al., 2017).   Some special classes of games in which an exact NE can be computed in polynomial time are:
@@ 2.3 Aggregative games
Aggregative games were initially studied for one-dimensional strategy sets (each agent chooses a number) and one-dimensional aggregators. Under certain conditions, such games admit a potential function, [^2] Roughly, a potential function is a function of the strategy profile, whose value increases when any player makes an improving move. A weaker definition only requires the potential to increase when any player makes a best response. and thus always have an NE (Kukushkin, 1994; Dubey et al., 2006; Cornes and Harley, 2012; Jensen, 2018). There are also efficient algorithms for computing an NE when players have discrete strategy sets (Kearns and Mansour, 2002)
@@
Against this background, our contribution is identifying a natural and practical class of aggregative games, for which an exact NE can be computed efficiently.
In the following, different types of aggregative games are covered.
@@
Even though Nash’s original existence proof is not valid in general for discrete strategy sets, congestion games (CGs) always have a potential function, so every sequence of individual improvements must terminate in a PNE. However, in general, this sequence might be exponentially long. Specifically, computing a PNE in general CGs is PLS-complete, which means that it probably cannot be done by a polytime algorithm (Fabrikant et al., 2004).333 PLS is similar to PPAD in that both contain problems in which a solution is guaranteed to exist — both are subsets of TFNP. It is believed that none of them contains the other, and that none of them is contained in P.
@@
  • Nonatomic CGs: CGs in which there is a continuum of players, where each player’s effect on the outcome is negligible. Finding a PNE in a nonatomic CG can be reduced to solving a convex optimization problem, which can be done in weakly polynomial time. In the special case of a network nonatomic CG with Lipschitz-continuous utility functions, there is an FPTAS (an algorithm computing an $\epsilon$-approximate PNE in time strongly-polynomial in the problem size and $1/\epsilon$). (Fabrikant et al., 2004).
@@
  • For symmetric (non-network) CGs, there is an FPTAS based on simple greedy dynamics (Chien and Sinclair, 2011).
  • For general CGs, when the delay functions are linear or polynomials of bounded degree, there are simple greedy dynamics that obtain a constant-factor approximate PNE (Caragiannis et al., 2015).
@@
The need to study NEs in budget-aggregation games is also reflected by the fact that many fairness-motivated solution concepts in participatory budgeting and multiwinner voting (see, e.g., Lackner and Skowron, 2023) are related to NEs, especially when agents have approval preferences, i.e., linear utilities with binary valuations, also called dichotomous utilities (Bogomolnaia et al., 2005). For the uncapped setting of budget-aggregation, when projects can receive arbitrary amounts of the budget, Brandl et al. (2022) coined the term “decomposability” for distributions of the overall budget that admit a decomposition into distributions of the individual budgets where each agent spends her budget optimally, i.e., only on approved projects. The standard model in participatory budgeting assumes that projects have fixed costs. For such problems, common rules like the method of equal shares (Peters and Skowron, 2020; Peters et al., 2021) or Phragmén’s rule (Phragmén, 1899; Janson, 2016) rely on the same idea that agents successively allocate their budgets to fund approved projects. Most of these works take a “mechanism design perspective” where agents report their preferences, which are then aggregated into a collective outcome, instead of individual budget distributions. However, for the special case of multiwinner approval voting (where all projects have the same costs and correspond to candidates), Haret et al. (2024) recently defined a budgeting game, which is very similar to our budget-aggregation game. In their game, as in ours, each player has a fixed budget, and can choose a distribution of this budget among $m$ issues. But in contrast to our game, their issues are discrete; each issue is active if its total allotment is at least $1$ and inactive otherwise. The utility of a player is discrete too — it is the number of active issues that the player approves. They prove that their budgeting game always has an NE, but do not elaborate on its computational aspects. Moreover, returning NEs already ensures some degree of fairness as they satisfy individual fair share (Bogomolnaia et al., 2005; Aziz et al., 2020) for arbitrary preference models (see Proposition 3.4).
@@ 3 Model
In a budget-aggregation game, the budget (w.l.o.g., normalized to $1$) is equally divided among the agents, and each of them is allowed to distribute their share among the alternatives. The set of strategies available to each payer is thus $S_{i}=\{\delta_{i}\in[0,1]^{m}\text{ such that }\sum_{x\in A}\delta_{i,x}=1/n\}$, which is an (m1)(m-1)-dimensional simplex in $\mathbb{R}^{m}$.
@@
The strategy of agent $i$, denoted $\delta_{i}=(\delta_{i,x})_{x\in A}$, implies that agent $i$ allocates $\delta_{i,x}$ to alternative $x$; these strategies are called individual distributions. We denote the overall distribution by $\delta:=\sum_{i\in N}\delta_{i}$. We assume that agents care only about the overall distribution of the budget, which is why the domain of $u_{i}$ is $\Delta(A)$ instead of $S_{1}\times\cdots\times S_{n}$. Furthermore, let $S_{-i}:=S_{1}\times\dots\times S_{i-1}\times S_{i+1}\times\dots\times S_{n}$ be the space of reduced strategy profiles where we exclude agent $i$.
@@
The aggregated distribution $\delta^{*}=\sum_{i\in N}\delta^{*}_{i}$ is called an equilibrium distribution.
@@
We are interested in understanding the structure and computing NEs for various utility models. In the following sections, we will consider linear utilities (Section 4), Leontief utilities (Section 5), binary symmetric separable utilities (Section 6), and $l_{1}$ disutilities (Section 7).
@@ 4 Linear utilities
Linear utilities, also known as additive or von Neumann-Morgenstern utility functions, are the most common extension of valuations for a set of pure alternatives $A$ to preferences over $\Delta(A)$. When agents’ preferences satisfy four seemingly mild conditions, von Neumann and Morgenstern (1944) showed that they can be represented by linear utility functions. Formally, an agent $i$ has linear utilities if for all $\delta\in\Delta(A)$,
@@
with nonnegative valuations $(v_{i,x})_{x\in A}$ for alternatives. [^4] Note that agents do not have different valuations for individual distributions $\delta_{i,x}$, $\delta_{j,x}$ to the same alternative $x$ as they only care about the total amount $\delta_{x}$ allocated to $x$.
@@
For linear utilities, the set of Nash equilibria consists of all $(\delta_{i})_{i\in N}$ with $\operatorname{supp}(\delta_{i})\subseteq A^{\max}_{i}$ for all $i\in N$.

### v3

@@ Abstract
Budget aggregation deals with the social choice problem of distributing an exogenously given budget among a set of public projects, given agents’ preferences. Taking a game-theoretic perspective, we study budget-aggregation games where each agent has virtual decision power over some fraction of the budget. We investigate the structure and show efficient computability of Nash equilibria for various common preference models in this setting. In particular, we show that equilibria for Leontief utilities can be found in polynomial time, solving an open problem from Brandt et al. (2023), and give an explicit polynomial-time algorithm for computing equilibria for $\ell_{1}$ preferences.
@@ 1 Introduction
Participatory budgeting (Cabannes, 2004; Aziz and Shah, 2021) is a democratic process in which citizens are allowed to vote on how the funds of their city or state should be distributed. It is justified both by ethical arguments (people have a fundamental right to decide how their money is used) and by practical arguments (letting people decide on the budget strengthens cohesion in the community).
This approach can be taken to the extreme: instead of just letting citizens vote on which projects they approve (as is commonly done today), one can give each citizen their own (virtual) share of the budget (e.g. $1/n$ of the total budget, in case there are $n$ citizens with equal entitlements), and allow them to decide how to allocate that share among the potential projects. This can be an attractive approach, as it gives full power to the citizens and ensures some degree of fairness by splitting this power equally among them. Furthermore, each citizen is able to directly observe her influence on the outcome by means of the distribution of her individual budget.
However, this approach raises severe strategic problems. For example, if a citizen thinks that projects $x$ and $y$ should each receive $0.5$ of the total budget, but later observed that most other citizens allocated their shares to $x$, then she would have preferred to spend her entire budget on $y$, leaving her regretful since her budget was spent sub-optimally in hindsight. This bad user experience can be avoided if we have the possibility to coordinate individual budgets by advising each citizen in advance how to distribute their own budget such that for each citizen, the advised distribution is optimal, given that all other citizens follow the provided advice.
@@
We propose a new class of normal-form games, called budget-aggregation games, which model (equal-resource) public funding situations. We position these games within the broader framework of normal-form games and survey existing results on the computation of NEs. On top of that, our paper analyzes NEs for the most common utility functions that have been considered in participatory budgeting and related fields. We prove theorems on the structure of NEs and develop algorithms to efficiently compute them. We focus on the following utility models:
@@
  • For Leontief utilities, we show that an NE can be computed in polynomial time using a variant of the ellipsoid method, solving an open problem from Brandt et al. (2023) (Section 5).
@@
Finally, we extend the basic model by allowing agents to have different weights. We show that, for most of the above utility models (namely linear, Leontief and $\ell_{1}$), an NE can still be computed in pseudo-polynomial time (Section 8).
@@ 2 Related Work
An important generalization of Nash’s theorem involves the notion of a concave game (Rosen, 1965). A concave game is defined by, for each player $i$, a convex set $(S_{i})_{i\in N}$ that represents the possible strategies for $i$, and a utility function $u_{i}$ that maps each strategy profile to a real number; $u_{i}$ can be any function that is continuous in all strategies, and concave in the strategy of ii. [^1] The definition in Rosen (1965) is even more general and allows coupled constraints, that is, the set of strategies available to each agent may depend on the choices made by other agents. We do not need this generalization here. Nash’s game is a special case in which every $S_{i}$ is a standard simplex in some Euclidean space (representing the set of all lotteries over pure actions), and every $u_{i}$ is a linear combination of products of $n$ terms. Our budget-aggregation game is a special case in which every $S_{i}$ is (still) a simplex, but the utility functions are substantially different, due to the interpretation as a distribution of exogenously given budget rather than a lottery.
2.1 Computing an equilibrium is hard
Like most fixed-point-based proofs, the ones by Nash (1950, 1951) and Rosen (1965) do not yield efficient algorithms for finding an equilibrium. Moreover, there are 3-player games with integer payoffs with a unique (mixed) NE in which all probabilities are irrational numbers (Nash, 1951; Bilò and Mavronicolas, 2014), and 4-player games with a unique NE in which all probabilities are irradical numbers (cannot be expressed using integer roots) (Orzech and Rinard, 2025). This shows that no finite algorithm can compute an exact NE in general games.
Even computing an $\epsilon$-approximate NE probably cannot be done in polynomial time, as it is PPAD-hard even for two players. This was first proven for $\epsilon$ that shrinks exponentially in the number of actions (Chen and Deng, 2006), then for $\epsilon$ that shrinks polynomially (Chen et al., 2006), and later even for some constant $\epsilon>0$ (Rubinstein, 2015). Computing an equilibrium is PPAD-complete, also for some other classes of concave games (Papadimitriou et al., 2023). Many natural decision problems regarding the existence of a NE with certain properties are NP-hard for two-player games, and $\exists\mathbb{R}$-hard for multi-player games (Bilò and Mavronicolas, 2014).
However, there are other classes of games for which an NE can be computed efficiently. We survey some of them below.
@@ 2.2 Nash games
For the original game studied by Nash (where the strategy set of each agent is the set of lotteries over a finite set of actions), many practical algorithms for computing an NE were designed, such as the ones by Lemke and Howson (1964) and Porter et al. (2008), but without a polynomial time guarantee.
Polynomial-time algorithms for $\epsilon$-approximate NE (where no agent can increase her utility by more than $\epsilon$ by deviating, where the utilities are normalized to $[0,1]$) are known for $\epsilon\approx 0.34$ for two players (Tsaknakis and Spirakis, 2007), $\epsilon\approx 0.5$ for $n$-player polymatrix games, and $\epsilon>0.6$ for general $n$-player games with $n\geq 3$ (Deligkas et al., 2017).   Some special classes of games in which an exact NE can be computed in polynomial time are:
@@ 2.3 Aggregative games
Aggregative games were initially studied for one-dimensional strategy sets (each agent chooses a number) and one-dimensional aggregators. Under certain conditions, such games admit a potential function, [^2] Roughly, a potential function is a function of the strategy profile, whose value increases when any player makes an improving move. A weaker definition only requires the potential to increase when any player makes a best response. and thus always have an NE (Kukushkin, 1994; Dubey et al., 2006; Cornes and Harley, 2012; Jensen, 2018). There are also efficient algorithms for computing an NE when players have discrete strategy sets (Kearns and Mansour, 2002).
@@
Against this background, our contribution is identifying a natural and practical class of aggregative games, for which an exact NE can be computed efficiently. In the following, different types of aggregative games are covered.
@@
Even though Nash’s original existence proof is not valid in general for discrete strategy sets, congestion games (CGs) always have a potential function, so every sequence of individual improvements must terminate in a PNE. However, in general, this sequence might be exponentially long. Specifically, computing a PNE in general CGs is PLS-complete, which means that it probably cannot be done by a polynomial-time algorithm (Fabrikant et al., 2004). [^3] PLS is similar to PPAD in that both contain problems in which a solution is guaranteed to exist — both are subsets of TFNP. It is believed that none of them contains the other, and that none of them is contained in P.
@@
  • Nonatomic CGs: CGs in which there is a continuum of players, where each player’s effect on the outcome is negligible (in contrast to our setting, where each player has a substantial effect on the outcome). Finding a PNE in a nonatomic CG can be reduced to solving a convex optimization problem, which can be done in weakly polynomial time. In the special case of a network nonatomic CG with Lipschitz-continuous utility functions, there is an FPTAS (an algorithm computing an $\epsilon$-approximate PNE in time strongly-polynomial in the problem size and $1/\epsilon$). (Fabrikant et al., 2004).
@@
  • For symmetric (non-network) CGs, an FPTAS based on simple greedy dynamics (Chien and Sinclair, 2011).
  • For general CGs, when the delay functions are polynomials of bounded degree, there are dynamics based on clever sequences of greedy steps that obtain a constant-factor approximate PNE (Caragiannis et al., 2015).
The CG class most closely related to our budget-aggregation game is the class of splittable (atomic) CGs (Orda et al., 2002; Roughgarden and Schoppmann, 2015), in which each agent controls a divisible load and can split it arbitrarily among different paths in a network. The players’ utilities are decreasing functions of the total load allocated to edges that they use: $u_{i}(\delta)=-\sum_{x\in A}\delta_{i,x}\cdot c_{i,x}(\delta_{x})$, where $c_{i,x}$ is the congestion function of edge $x$ for player $i$. An important special case is when the congestion functions are affine, $c_{i,x}(\delta_{x})=a_{i,x}\cdot\delta_{x}+b_{i,x}$. Klimm and Warode (2025) prove that, even with affine congestion functions, computing a PNE is PPAD-complete. Still, polynomial-time algorithms are known for some special cases. If the congestion functions are affine and player-independent, an approximate PNE can be computed by convex programming for general graphs (Cominetti et al., 2009); for specific classes of graphs called ”well-designed”, an exact PNE can be computed by a combinatorial algorithm (Huang, 2013). When strategy sets are singletons — the strategy set of each player is a set of edges, rather than a set of paths from a source to a target — there is a polynomial-time algorithm for affine player-specific congestion functions (Harks and Timmermans, 2022); for convex player-independent congestion functions, there are algorithms that are polynomial when either the number of players or the number of edges is fixed (Bhaskar and Lolakapuri, 2018).
Note that the utility functions are different than our linear utility functions in that the (negative) utility that a player derives from an edge is multiplied by the player’s own contribution to that edge. In other words, the dependence of $u_{i}(\delta)$ on $\delta_{i,x}$ is quadratic, as $\delta_{i,x}$ both multiplies $c_{i,x}$ and contributes to $\delta_{x}$.
@@
The need to study NEs in budget-aggregation games is also reflected by the fact that many fairness-motivated solution concepts in participatory budgeting and multiwinner voting (see, e.g., Lackner and Skowron, 2023) are related to NEs, especially when agents have approval preferences, i.e., linear utilities with binary valuations, also called dichotomous utilities (Bogomolnaia et al., 2005). For the uncapped setting of budget-aggregation, when projects can receive arbitrary amounts of the budget, Brandl et al. (2022) coined the term “decomposability” for distributions of the overall budget that admit a decomposition into distributions of the individual budgets where each agent spends her budget optimally, i.e., only on approved projects. The standard model in participatory budgeting assumes that projects have fixed costs. For such problems, common rules like the method of equal shares (Peters and Skowron, 2020; Peters et al., 2021) or Phragmén’s rule (Phragmén, 1899; Janson, 2016) rely on the same idea that agents successively allocate their budgets to fund approved projects. Most of these works take a “mechanism design perspective” where agents report their preferences, which are then aggregated into a collective outcome, instead of individual budget distributions. A notable exception is donor coordination (see, e.g., Freeman et al., 2019) where the budget is initially owned by the agents. Brandt et al. (2023, 2025a) proved that, with Leontief utilities, the unique equilibrium distribution is the unique distribution that maximizes the product of agents’ utilities where agents are weighted by their contributions. For the special case of multiwinner approval voting (where all projects have the same costs and correspond to candidates), Haret et al. (2024) recently defined a budgeting game, which is very similar to our budget-aggregation game. In their game, as in ours, each player has a fixed budget, and can choose a distribution of this budget among $m$ issues. But in contrast to our game, their issues are discrete; each issue is active if its total allotment is at least $1$ and inactive otherwise. The utility of a player is the number of active issues that the player approves. They prove that their budgeting game always has an NE, but do not elaborate on its computational aspects. Moreover, returning NEs already ensures some degree of fairness as they satisfy individual fair share (Bogomolnaia et al., 2005; Aziz et al., 2020) for arbitrary preference models (see Proposition 3.4).
@@ 3 Model
In a budget-aggregation game, the budget (w.l.o.g., normalized to $1$) is equally divided among the agents, and each of them is allowed to distribute their share among the alternatives. The set of strategies available to each player is thus $S_{i}=\{\delta_{i}\in[0,1]^{m}\text{ such that }\sum_{x\in A}\delta_{i,x}=1/n\}$, which is an (m1)(m-1)-dimensional simplex in $\mathbb{R}^{m}$.
@@
The strategy of agent $i$, denoted $\delta_{i}=(\delta_{i,x})_{x\in A}$, implies that agent $i$ allocates $\delta_{i,x}$ to alternative $x$; these strategies are called individual distributions. We denote the overall distribution by $\delta:=\sum_{i\in N}\delta_{i}$. Note that agents only care about the overall distribution of the budget, which is why the domain of $u_{i}$ is $\Delta(A)$ instead of $S_{1}\times\cdots\times S_{n}$. Furthermore, let $S_{-i}:=S_{1}\times\dots\times S_{i-1}\times S_{i+1}\times\dots\times S_{n}$ be the space of reduced strategy profiles where we exclude agent $i$.
@@
The sum $\delta^{*}=\sum_{i\in N}\delta^{*}_{i}$ is called an equilibrium distribution.
@@
We are interested in understanding the structure and computing NEs for various utility models. In the following sections, we will consider linear utilities (Section 4), Leontief utilities (Section 5), binary symmetric separable utilities (Section 6), and $\ell_{1}$ disutilities (Section 7).
@@ 4 Linear utilities
Linear utilities, also known as additive or von Neumann-Morgenstern utility functions, are the most common extension of valuations for a set of pure alternatives $A$ to preferences over $\Delta(A)$. When agents’ preferences satisfy four seemingly mild conditions, von Neumann and Morgenstern (1944) showed that they can be represented by linear utility functions. Formally, $i$ has linear utilities if for all $\delta\in\Delta(A)$,
@@
with nonnegative valuations $(v_{i,x})_{x\in A}$ for alternatives. [^4] Agents do not have different valuations for individual distributions $\delta_{i,x}$, $\delta_{j,x}$ to the same alternative $x$ as they only care about the total amount $\delta_{x}$ allocated to $x$.
@@
For linear utilities, the set of NE consists of all $(\delta_{i})_{i\in N}$ with $\operatorname{supp}(\delta_{i})\subseteq A^{\max}_{i}$ for all $i\in N$.  [^5] If we interpret the overall distribution as a probability distribution, Proposition 4.1 implies that a distribution is in equilibrium if and only if it is the outcome of a random dictatorship, where a player is selected uniformly at random, and then selects one of his best outcomes. Random dictatorship is sometimes used as a benchmark of fairness; see, e.g., Bogomolnaia et al. (2005).